The square root of 12 -5x= 5x
Answer by nyc_function(2741)

sqrt(12 - 5x} = 5x
Square both sides.
[sqrt{12 - 5x}]^2 = (5x)^2
12 - 5x = 25x^2
Put everything on one side and equate to zero.
25x^2 + 5x - 12 = 0
You now have a quadratic equation.
Can you take it from here?